# IRAD: Implicit Representation-driven Image Resampling against Adversarial Attacks

### v1

@@ 6.1 Comparing with SOTA Methods
IRAD+DiffPure (t=20) 93.42 74.05 83.73 27.70
❹ Following the relationship between DISCO and IRAD as discussed in Sec. 5, the substantial advantages of IRAD over DISCO underscore the effectiveness of the proposed SampleNet.
Adaptive adversary scenario. We further study the performance of all methods under a more challenging scenario where the attack is aware of both the defense methods and classification models and employs BPDA (Athalye et al., 2018) as the attack. As shown in Table 3, we have the following observations: ❶ RAs of all methods reduce under the BPDA. For example, DISCO gets 85.63% RA against AutoAttack but only achieves 22.60% RA under BPDA on the CIFAR10. The RA of our method reduces from 89.72% to 74.32% on CIFAR10. ❷ Our method achieves the highest RA among all compared methods on all three datasets, with the exception of DiffPure in Cifar10. Although the RA of our method reduces, the relative improvements over other methods become much larger. DiffPure’s high performance comes at the expense of significant computational resources. Therefore, we suggest a hybrid approach that capitalizes on the strengths of both DiffPure and IRAD, as detailed in the following section. ❸ Our method still achieves the highest average accuracy when compared to all other methods, except for DiffPure in CIFAR-10, highlighting the advantages of IRAD in achieving a favorable trade-off between SA and RA under the adaptive adversary scenario. AutoAttack-based Adaptive adversary scenario. We also explored a tough defense scenario where attackers have full knowledge of defense methods and models. In Table 5, we found that DISCO and our original IRAD didn’t improve robustness with zero RA. DiffPure maintains high SA and RA under adaptive settings but takes 132 ms per image, which is time-consuming. To speed up DiffPure, we can use fewer time steps, but this reduces RA significantly. When we combined IRAD with DiffPure (See Sec. 5), our IRAD+DiffPure (t=20) method achieved comparable SA and RA with DiffPure, and it’s about five times faster.
@@ 6.2 Generalization across DNNs, Datasets, and Other Attacks
CIFAR10 89.72 68.41
CIFAR100 89.06 71.76
Generalization against other attacks We further use IRAD trained via PGD-based adversarial examples against 10 attacking methods and compare it with the SOTA method DISCO. As shown in Fig. 4, we have the following observations: ❶ In the CIFAR10 and ImageNet datasets, IRAD achieves similar RA across all attack methods, which demonstrates the high generalization of IRAD across diverse attacks and even unknown attacks. Regarding the CIFAR100 dataset, IRAD presents slightly lower RA under MiFGSM than other attacks. ❷ With the exception of the APdg attack on CIFAR100, where DISCO marginally outperforms IRAD, our approach consistently achieves significantly higher robust accuracies (RAs) across all other attack scenarios, underscoring its advantages.
@@ 6.3 Ablation Study
Sampling strategies comparison. We replace the SampleNet of IRAD with two naive sampling strategies introduced in Sec. 3.2 and evaluate the performance on the CIFAR10 dataset via AutoAttack ($\epsilon_{\infty}=8/255$). As shown in Table 8, we see that SampleNet can achieve the highest SA, RA, and average accuracy, which demonstrates the effectiveness of our SampleNet.
@@ 7 Conclusion
We have identified a novel adversarial defense solution, i.e., image resampling, which can break the adversarial textures while maintaining the main semantic information in the input image. We provided a general formulation for the image resampling-based adversarial defense and designed several naive defensive resampling methods. We further studied the effectiveness and limitations of these naive methods. To fill the limitations, we proposed the implicit continuous representation-driven image resampling method by building the implicit representation and designing a SampleNet that can predict coordinate shifting magnitudes for all pixels according to different inputs. The experiments have demonstrated the advantages of our method over existing methods. In the future, this method could be combined with the other two defensive methods, i.e., denoising and adversarial training, for constructing much higher robust models.
Appendix A Supplementary Material
@@ A.1 Experiment Settings in Detail
For SampleNet training, we utilize cross-entropy loss, considering the prediction results of both clean and attacked images as the loss function. For optimization, we employ the Adam optimizer with a learning rate of 2e-4 and betas set to (0, 0.9). The training of SampleNet is performed with a batch size of 400 for CIFAR10 and CIFAR100, and with a batch size of 8 for ImageNet.
@@ A.2 Comparing with More SOTA Methods on Cifar10 under AutoAttack-based Adaptive adversary scenario
In this section, we present additional results from various methods comparison on Cifar10 under AutoAttack-based adaptive adversary scenario. As presented in Table 9, we present the baseline results from RobustBench, and compare them with the result we provided on Table 5.
From Table 9, we can conclude that IRAD+DiffPure method we proposed has advantages over existing SOTA methods on RobustBench.
@@
Defense
SA RA Avg.
Classifier
w.o. Defense
94.77 0 47.39
WideResNet-28-10
Robust_Peng Peng et al. (2023)
93.27 71.07 82.17
RaWideResNet-70-16
Better_Wang Wang et al. (2023)
93.25 70.69 81.97
WideResNet-70-16
Better_Wang Wang et al. (2023)
92.44 67.31 79.88
WideResNet-28-10
Fixing_Rebuffi Rebuffi et al. (2021)
92.23 66.58 79.41
WideResNet-70-16
Improving_Gowal Gowal et al. (2021)
88.74 66.11 77.43
WideResNet-70-16
Uncovering_Gowal Gowal et al. (2020)
91.1 65.88 78.49
WideResNet-70-16
Revisiting_HuangHuang et al. (2022)
91.58 65.79 78.69
WideResNet-A4
Fixing_Rebuffi Rebuffi et al. (2021)
88.5 64.64 76.57
WideResNet-106-16
Stable_Kang Kang et al. (2021)
93.73 71.28 82.51
WideResNet-70-16, Neural ODE block
Fixing_Rebuffi Rebuffi et al. (2021)
88.54 64.25 76.40
WideResNet-70-16
Improving_Gowal Gowal et al. (2021)
87.5 63.44 75.47
WideResNet-28-10
Robustness_Pang Pang et al. (2022)
89.01 63.35 76.18
WideResNet-70-16
Helper_RadeRade & Moosavi-Dezfooli (2021)
91.47 62.83 77.15
WideResNet-34-10
Robust_Sehwag Sehwag et al. (2021)
87.3 62.79 75.05
ResNest152
Uncovering_Gowal Gowal et al. (2020)
89.48 62.8 76.14
WideResNet-28-10
Exploring_Huang Huang et al. (2021a)
91.23 62.54 76.89
WideResNet-34-R
Exploring_Huang Huang et al. (2021a)
90.56 61.56 76.06
WideResNet-34-R
Parameterizing_Dai Dai et al. (2022)
87.02 61.55 74.29
WideResNet-28-10-PSSiLU
Robustness_Pang Pang et al. (2022)
88.61 61.04 74.83
WideResNet-28-10
Helper_RadeRade & Moosavi-Dezfooli (2021)
88.16 60.97 74.57
WideResNet-28-10
Fixing_Rebuffi Rebuffi et al. (2021)
87.33 60.75 74.04
WideResNet-28-10
Wider_Wu Wu et al. (2021)
87.67 60.65 74.16
WideResNet-34-15
Improving_Sridhar Sridhar et al. (2022)
86.53 60.41 73.47
WideResNet-34-15
Robust_Sehwag Sehwag et al. (2021)
86.68 60.27 73.48
WideResNet-34-10
Adversarial_Wu Wu et al. (2020)
88.25 60.04 74.15
WideResNet-28-10
Improving_Sridhar Sridhar et al. (2022)
89.46 59.66 74.56
WideResNet-28-10
Geometry_Zhang Zhang et al. (2020b)
89.36 59.64 74.50
WideResNet-28-10
Unlabeled_Carmon Carmon et al. (2019)
89.69 59.53 74.61
WideResNet-28-10
Improving_Gowal Gowal et al. (2021)
87.35 58.63 72.99
PreActResNet-18
Towards_Addepalli Addepalli et al. (2021)
85.32 58.04 71.68
WideResNet-34-10
Efficient_Addepalli Addepalli et al. (2022)
88.71 57.81 73.26
WideResNet-34-10
Ltd_Chen Chen & Lee (2021)
86.03 57.71 71.87
WideResNet-34-20
Helper_RadeRade & Moosavi-Dezfooli (2021)
89.02 57.67 73.35
PreActResNet-18
Adversarial_Jia Jia et al. (2022)
85.66 57.61 71.64
WideResNet-70-16
Light_Debenedetti Debenedetti et al. (2022)
91.73 57.58 74.66
XCiT-L12
Light_Debenedetti Debenedetti et al. (2022)
91.3 57.27 74.29
XCiT-M12
Uncovering_Gowal Gowal et al. (2020)
85.29 57.2 71.25
WideResNet-70-16
Hydra_SehwagSehwag et al. (2020)
88.98 57.14 73.06
WideResNet-28-10
Helper_RadeRade & Moosavi-Dezfooli (2021)
86.86 57.09 71.98
PreActResNet-18
Ltd_Chen Chen & Lee (2021)
85.21 56.94 71.08
WideResNet-34-10
Uncovering_Gowal Gowal et al. (2020)
85.64 56.86 71.25
WideResNet-34-20
Fixing_Rebuffi Rebuffi et al. (2021)
83.53 56.66 70.10
PreActResNet-18
Improving_Wang Wang et al. (2020)
87.5 56.29 71.90
WideResNet-28-10
Adversarial_Jia Jia et al. (2022)
84.98 56.26 70.62
WideResNet-34-10
Adversarial_Wu Wu et al. (2020)
85.36 56.17 70.77
WideResNet-34-10
Light_Debenedetti Debenedetti et al. (2022)
90.06 56.14 73.10
XCiT-S12
Labels_Uesato Uesato et al. (2019)
86.46 56.03 71.25
WideResNet-28-10
Robust_Sehwag Sehwag et al. (2021)
84.59 55.54 70.07
ResNet-18
Using_Hendrycks Hendrycks et al. (2019)
87.11 54.92 71.02
WideResNet-28-10
Bag_Pang Pang et al. (2020a)
86.43 54.39 70.41
WideResNet-34-20
Boosting_Pang Pang et al. (2020b)
85.14 53.74 69.44
WideResNet-34-20
Learnable_Cui Cui et al. (2021)
88.7 53.57 71.14
WideResNet-34-20
Attacks_Zhang Zhang et al. (2020a)
84.52 53.51 69.02
WideResNet-34-10
Overfitting_RiceRice et al. (2020)
85.34 53.42 69.38
WideResNet-34-20
Self_Huang Huang et al. (2020)
83.48 53.34 68.41
WideResNet-34-10
Theoretically_ZhangZhang et al. (2019b)
84.92 53.08 69.00
WideResNet-34-10
Learnable_Cui Cui et al. (2021)
88.22 52.86 70.54
WideResNet-34-10
Adversarial_Qin Qin et al. (2019)
86.28 52.84 69.56
WideResNet-40-8
Efficient_Addepalli Addepalli et al. (2022)
85.71 52.48 69.10
ResNet-18
Adversarial_Chen Chen et al. (2020)
86.04 51.56 68.80
ResNet-50
Efficient_Chen Chen et al. (2022)
85.32 51.12 68.22
WideResNet-34-10
Towards_Addepalli Addepalli et al. (2021)
80.24 51.06 65.65
ResNet-18
Improving_Sitawarin Sitawarin et al. (2020)
86.84 50.72 68.78
WideResNet-34-10
Robustness Engstrom et al. (2019)
87.03 49.25 68.14
ResNet-50
Harnessing_Kumari Kumari et al. (2019)
87.8 49.12 68.46
WideResNet-34-10
Metric_Mao Mao et al. (2019)
86.21 47.41 66.81
WideResNet-34-10
You_Zhang Zhang et al. (2019a)
87.2 44.83 66.02
WideResNet-34-10
Towards_Madry Madry et al. (2017)
87.14 44.04 65.59
WideResNet-34-10
Understanding_AndriushchenkoAndriushchenko & Flammarion (2020)
79.84 43.93 61.89
PreActResNet-18
Rethinking_Pang Pang et al. (2019)
80.89 43.48 62.19
ResNet-32
Fast_Wong Wong et al. (2020)
83.34 43.21 63.28
PreActResNet-18
Adversarial_Shafahi Shafahi et al. (2019)
86.11 41.47 63.79
WideResNet-34-10
Mma_Ding Ding et al. (2018)
84.36 41.44 62.90
WideResNet-28-4
Tunable_KunduKundu et al. (2020)
87.32 40.41 63.87
ResNet-18
Controlling_Atzmon Atzmon et al. (2019)
81.3 40.22 60.76
ResNet-18
Robustness_Moosavi Moosavi-Dezfooli et al. (2019)
83.11 38.5 60.81
ResNet-18
Defense_Zhang Zhang & Wang (2019)
89.98 36.64 63.31
WideResNet-28-10
Adversarial_Zhang Zhang & Xu
90.25 36.45 63.35
WideResNet-28-10
Adversarial_JangJang et al. (2019)
78.91 34.95 56.93
ResNet-20
Sensible_Kim Kim & Wang (2020)
91.51 34.22 62.87
WideResNet-34-10
Adversarial_ZhangZhang & Xu
44.73 32.64 38.69
5-layer-CNN
Bilateral_Wang Wang & Zhang (2019)
92.8 29.35 61.08
WideResNet-28-10
Enhancing_Xiao Xiao et al. (2019)
79.28 18.5 48.89
DenseNet-121
Manifold_JinJin & Rinard (2020)
90.84 1.35 46.10
ResNet-18
Adversarial_Mustafa Mustafa et al. (2019)
89.16 0.28 44.72
ResNet-110
Jacobian_Chan Chan et al. (2019)
93.79 0.26 47.03
WideResNet-34-10
Clustr_Alfarra Alfarra et al. (2020)
91.03 0 45.52
WideResNet-28-10
IRAD+DiffPure (t=20)
93.42 74.05 83.73
WideResNet-28-10

### v3

@@ 6.1 Comparing with SOTA Methods
IRAD+DiffPure (t=20) 93.42 74.05 83.74 27.70
❹ Following the relationship between DISCO and IRAD as discussed in Sec. 5, the substantial advantages of IRAD over DISCO underscore the effectiveness of the proposed SampleNet. Adaptive adversary scenario. We further study the performance of all methods under a more challenging scenario where the attack is aware of both the defense methods and classification models and employs BPDA (Athalye et al., 2018) as the attack. As shown in Table 3, we have the following observations: ❶ RAs of all methods reduce under the BPDA. For example, DISCO gets 85.63% RA against AutoAttack but only achieves 22.60% RA under BPDA on CIFAR10. The RA of our method reduces from 89.72% to 74.32% on CIFAR10. ❷ Our method achieves the highest RA among all compared methods on all three datasets, with the exception of DiffPure in Cifar10. Although the RA of our method reduces, the relative improvements over other methods become much larger. DiffPure’s high performance comes at the expense of significant computational resources. Therefore, we suggest a hybrid approach that combines the strengths of DiffPure and IRAD, as detailed in the following section. ❸ Our method still achieves the highest average accuracy when compared to all other methods, except for DiffPure in CIFAR-10, highlighting the advantages of IRAD in achieving a favorable trade-off between SA and RA under the adaptive adversary scenario. AutoAttack-based Adaptive adversary scenario. We also explored a tough defense scenario where attackers have full knowledge of defense methods and models. In Table 5, we found that DISCO and our original IRAD didn’t improve robustness with zero RA. DiffPure maintains high SA and RA under adaptive settings but takes 132 ms per image, which is time-consuming. To speed up DiffPure, we can use fewer time steps, but this reduces RA significantly. When we combined IRAD with DiffPure (See Sec. 5), our IRAD+DiffPure (t=20) method achieved comparable SA and RA with DiffPure, and it’s about five times faster. More experiments comparing different steps in the integration of IRAD and DiffPure are in Appendix A.6.
@@ 6.2 Generalization across DNNs, Datasets, and Other Attacks
CIFAR10 89.72 67.14
CIFAR100 88.68 72.49
Generalization against other attacks. We further use IRAD trained via PGD-based adversarial examples against 10 attacking methods and compare it with the SOTA method DISCO. As shown in Fig. 4, we have the following observations: ❶ In the CIFAR10 and ImageNet datasets, IRAD achieves similar RA across all attack methods, which demonstrates the high generalization of IRAD across diverse attacks and even unknown attacks. Regarding the CIFAR100 dataset, IRAD presents slightly lower RA under MiFGSM than other attacks. ❷ IRAD consistently achieves significantly higher robust accuracies (RAs) across all attack scenarios, underscoring its advantages. Detailed results and additional experiments on more attacks can be found in Appendix A.3 and A.4.
@@ 6.3 Ablation Study
Sampling strategies comparison. We replace the SampleNet of IRAD with two naive sampling strategies introduced in Sec. 3.2 and evaluate the performance on the CIFAR10 dataset via AutoAttack ($\epsilon_{\infty}=8/255$). As shown in Table 8, we see that SampleNet can achieve the highest SA, RA, and average accuracy, which demonstrates the effectiveness of our SampleNet. Additional results regarding the ablation study and the effectiveness analysis can be found in the Appendix.
@@ 7 Conclusion
We have identified a novel adversarial defense solution, i.e., image resampling, which can break the adversarial textures while maintaining the main semantic information in the input image. We provided a general formulation for the image resampling-based adversarial defense and designed several naive defensive resampling methods. We further studied the effectiveness and limitations of these naive methods. To fill the limitations, we proposed the implicit continuous representation-driven image resampling method by building the implicit representation and designing a SampleNet that can predict coordinate shifting magnitudes for all pixels according to different inputs. The experiments have demonstrated the advantages of our method over all existing methods. In the future, this method could be combined with the other two defensive methods, i.e., denoising and adversarial training, for constructing much higher robust models.
Acknowledgment
This research is supported by the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-GC-2023-008), and Career Development Fund (CDF) of the Agency for Science, Technology and Research (A*STAR) (No.: C233312028). Xiaofeng Cao is supported by the National Natural Science Foundation of China (Grant Number: 62206108). The research is also supported by the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and Cyber Security Agency of Singapore.
Appendix A Appendix
@@ A.1 Experiment Settings in Detail
For SampleNet training, we utilize cross-entropy loss, considering the prediction results of both clean and attacked images as the loss function. For optimization, we employ the Adam optimizer with a learning rate of 2e-4 for CIFAR10 and ImageNet, 1e-3 for CIFAR100, and betas set to (0, 0.9). The training of SampleNet is conducted with a batch size of 400 for CIFAR10, 200 for CIFAR100, and 8 for ImageNet.
@@ A.2 Comparing with More SOTA Methods on Cifar10 under AutoAttack-based Adaptive adversary scenario
In this section, we present additional results from various methods comparison on Cifar10 under AutoAttack-based adaptive adversary scenario. As presented in Table LABEL:tab:SOTA_Cifar10_Adaptive_AutoAttack, we present the baseline results from RobustBench, and compare them with the result we provided on Table 5 and Table 16.
From Table LABEL:tab:SOTA_Cifar10_Adaptive_AutoAttack, we can conclude that IRAD+DiffPure method we proposed has advantages over existing SOTA methods on RobustBench.
@@
Defense SA RA Avg. Classifier
w.o. Defense 94.77 0 47.39 WideResNet-28-10
Robust_Peng (Peng et al., 2023) 93.27 71.07 82.17 RaWideResNet-70-16
Better_Wang (Wang et al., 2023) 93.25 70.69 81.97 WideResNet-70-16
Better_Wang (Wang et al., 2023) 92.44 67.31 79.88 WideResNet-28-10
Fixing_Rebuffi (Rebuffi et al., 2021) 92.23 66.58 79.41 WideResNet-70-16
Improving_Gowal (Gowal et al., 2021) 88.74 66.11 77.43 WideResNet-70-16
Uncovering_Gowal (Gowal et al., 2020) 91.1 65.88 78.49 WideResNet-70-16
Revisiting_Huang (Huang et al., 2022) 91.58 65.79 78.69 WideResNet-A4
Fixing_Rebuffi (Rebuffi et al., 2021) 88.5 64.64 76.57 WideResNet-106-16
Stable_Kang (Kang et al., 2021) 93.73 71.28 82.51 WideResNet-70-16, Neural ODE block
Fixing_Rebuffi (Rebuffi et al., 2021) 88.54 64.25 76.40 WideResNet-70-16
Improving_Gowal (Gowal et al., 2021) 87.5 63.44 75.47 WideResNet-28-10
Robustness_Pang (Pang et al., 2022) 89.01 63.35 76.18 WideResNet-70-16
Helper_Rade(Rade & Moosavi-Dezfooli, 2021) 91.47 62.83 77.15 WideResNet-34-10
Robust_Sehwag (Sehwag et al., 2021) 87.3 62.79 75.05 ResNest152
Uncovering_Gowal (Gowal et al., 2020) 89.48 62.8 76.14 WideResNet-28-10
Exploring_Huang (Huang et al., 2021a) 91.23 62.54 76.89 WideResNet-34-R
Exploring_Huang (Huang et al., 2021a) 90.56 61.56 76.06 WideResNet-34-R
Parameterizing_Dai (Dai et al., 2022) 87.02 61.55 74.29 WideResNet-28-10-PSSiLU
Robustness_Pang (Pang et al., 2022) 88.61 61.04 74.83 WideResNet-28-10
Helper_Rade(Rade & Moosavi-Dezfooli, 2021) 88.16 60.97 74.57 WideResNet-28-10
Fixing_Rebuffi (Rebuffi et al., 2021) 87.33 60.75 74.04 WideResNet-28-10
Wider_Wu (Wu et al., 2021) 87.67 60.65 74.16 WideResNet-34-15
Improving_Sridhar (Sridhar et al., 2022) 86.53 60.41 73.47 WideResNet-34-15
Robust_Sehwag (Sehwag et al., 2021) 86.68 60.27 73.48 WideResNet-34-10
Adversarial_Wu (Wu et al., 2020) 88.25 60.04 74.15 WideResNet-28-10
Improving_Sridhar (Sridhar et al., 2022) 89.46 59.66 74.56 WideResNet-28-10
Geometry_Zhang (Zhang et al., 2020c) 89.36 59.64 74.50 WideResNet-28-10
Unlabeled_Carmon (Carmon et al., 2019) 89.69 59.53 74.61 WideResNet-28-10
Improving_Gowal (Gowal et al., 2021) 87.35 58.63 72.99 PreActResNet-18
Towards_Addepalli (Addepalli et al., 2021) 85.32 58.04 71.68 WideResNet-34-10
Efficient_Addepalli (Addepalli et al., 2022) 88.71 57.81 73.26 WideResNet-34-10
Ltd_Chen (Chen & Lee, 2021) 86.03 57.71 71.87 WideResNet-34-20
Helper_Rade(Rade & Moosavi-Dezfooli, 2021) 89.02 57.67 73.35 PreActResNet-18
Adversarial_Jia (Jia et al., 2022a) 85.66 57.61 71.64 WideResNet-70-16
Light_Debenedetti (Debenedetti et al., 2022) 91.73 57.58 74.66 XCiT-L12
Light_Debenedetti (Debenedetti et al., 2022) 91.3 57.27 74.29 XCiT-M12
Uncovering_Gowal (Gowal et al., 2020) 85.29 57.2 71.25 WideResNet-70-16
Hydra_Sehwag(Sehwag et al., 2020) 88.98 57.14 73.06 WideResNet-28-10
Helper_Rade(Rade & Moosavi-Dezfooli, 2021) 86.86 57.09 71.98 PreActResNet-18
Ltd_Chen (Chen & Lee, 2021) 85.21 56.94 71.08 WideResNet-34-10
Uncovering_Gowal (Gowal et al., 2020) 85.64 56.86 71.25 WideResNet-34-20
Fixing_Rebuffi (Rebuffi et al., 2021) 83.53 56.66 70.10 PreActResNet-18
Improving_Wang (Wang et al., 2020) 87.5 56.29 71.90 WideResNet-28-10
Adversarial_Jia (Jia et al., 2022a) 84.98 56.26 70.62 WideResNet-34-10
Adversarial_Wu (Wu et al., 2020) 85.36 56.17 70.77 WideResNet-34-10
Light_Debenedetti (Debenedetti et al., 2022) 90.06 56.14 73.10 XCiT-S12
Labels_Uesato (Uesato et al., 2019) 86.46 56.03 71.25 WideResNet-28-10
Robust_Sehwag (Sehwag et al., 2021) 84.59 55.54 70.07 ResNet-18
Using_Hendrycks (Hendrycks et al., 2019) 87.11 54.92 71.02 WideResNet-28-10
Bag_Pang (Pang et al., 2020a) 86.43 54.39 70.41 WideResNet-34-20
Boosting_Pang (Pang et al., 2020b) 85.14 53.74 69.44 WideResNet-34-20
Learnable_Cui (Cui et al., 2021) 88.7 53.57 71.14 WideResNet-34-20
Attacks_Zhang (Zhang et al., 2020b) 84.52 53.51 69.02 WideResNet-34-10
Overfitting_Rice(Rice et al., 2020) 85.34 53.42 69.38 WideResNet-34-20
Self_Huang (Huang et al., 2020) 83.48 53.34 68.41 WideResNet-34-10
Theoretically_Zhang(Zhang et al., 2019b) 84.92 53.08 69.00 WideResNet-34-10
Learnable_Cui (Cui et al., 2021) 88.22 52.86 70.54 WideResNet-34-10
Adversarial_Qin (Qin et al., 2019) 86.28 52.84 69.56 WideResNet-40-8
Efficient_Addepalli (Addepalli et al., 2022) 85.71 52.48 69.10 ResNet-18
Adversarial_Chen (Chen et al., 2020) 86.04 51.56 68.80 ResNet-50
Efficient_Chen (Chen et al., 2022) 85.32 51.12 68.22 WideResNet-34-10
Towards_Addepalli (Addepalli et al., 2021) 80.24 51.06 65.65 ResNet-18
Improving_Sitawarin (Sitawarin et al., 2020) 86.84 50.72 68.78 WideResNet-34-10
Robustness (Engstrom et al., 2019) 87.03 49.25 68.14 ResNet-50
Harnessing_Kumari (Kumari et al., 2019) 87.8 49.12 68.46 WideResNet-34-10
Metric_Mao (Mao et al., 2019) 86.21 47.41 66.81 WideResNet-34-10
You_Zhang (Zhang et al., 2019a) 87.2 44.83 66.02 WideResNet-34-10
Towards_Madry (Madry et al., 2017) 87.14 44.04 65.59 WideResNet-34-10
Understanding_Andriushchenko(Andriushchenko & Flammarion, 2020) 79.84 43.93 61.89 PreActResNet-18
Rethinking_Pang (Pang et al., 2019) 80.89 43.48 62.19 ResNet-32
Fast_Wong (Wong et al., 2020) 83.34 43.21 63.28 PreActResNet-18
Adversarial_Shafahi (Shafahi et al., 2019) 86.11 41.47 63.79 WideResNet-34-10
Mma_Ding (Ding et al., 2018) 84.36 41.44 62.90 WideResNet-28-4
Tunable_Kundu(Kundu et al., 2020) 87.32 40.41 63.87 ResNet-18
Controlling_Atzmon (Atzmon et al., 2019) 81.3 40.22 60.76 ResNet-18
Robustness_Moosavi (Moosavi-Dezfooli et al., 2019) 83.11 38.5 60.81 ResNet-18
Defense_Zhang (Zhang & Wang, 2019) 89.98 36.64 63.31 WideResNet-28-10
Adversarial_Zhang (Zhang & Xu, ) 90.25 36.45 63.35 WideResNet-28-10
Adversarial_Jang(Jang et al., 2019) 78.91 34.95 56.93 ResNet-20
Sensible_Kim (Kim & Wang, 2020) 91.51 34.22 62.87 WideResNet-34-10
Adversarial_Zhang(Zhang & Xu, ) 44.73 32.64 38.69 5-layer-CNN
Bilateral_Wang (Wang & Zhang, 2019) 92.8 29.35 61.08 WideResNet-28-10
Enhancing_Xiao (Xiao et al., 2019) 79.28 18.5 48.89 DenseNet-121
Manifold_Jin(Jin & Rinard, 2020) 90.84 1.35 46.10 ResNet-18
Adversarial_Mustafa (Mustafa et al., 2019) 89.16 0.28 44.72 ResNet-110
Jacobian_Chan (Chan et al., 2019) 93.79 0.26 47.03 WideResNet-34-10
Clustr_Alfarra (Alfarra et al., 2020) 91.03 0 45.52 WideResNet-28-10
IRAD+DiffPure (t=20) 93.42 74.05 83.74 WideResNet-28-10
IRAD+DiffPure (t=25) 91.42 77.71 84.57 WideResNet-28-10
IRAD+DiffPure (t=30) 92.33 82.85 87.59 WideResNet-28-10
IRAD+DiffPure (t=35) 89.14 84.57 86.86 WideResNet-28-10
IRAD+DiffPure (t=40) 90.57 86.00 88.29 WideResNet-28-10